OpenStudy (ray10):

use partial derivatives to approximate the value of \[f(6.1,1.9) \] where \[f(x,y) = (x^{2}-y^{3}-1)^{\frac{ 1 }{ 3 }}\] at \[(6,2)\]

OpenStudy (anonymous):

partial derivative with respect to x or y

OpenStudy (anonymous):

or both?

OpenStudy (ray10):

The question doesn't state :S so I think both :)

OpenStudy (anonymous):

I'll do with respect to x first

OpenStudy (anonymous):

what is the partial derivative of f with respect to x? do you know how to do this?

OpenStudy (ray10):

sure thing :)

OpenStudy (ray10):

yeah I think I know :)

OpenStudy (anonymous):

okay so compute it and when you have this you simply plug in x=6 and y=2 the same for the partial with respect to y

OpenStudy (anonymous):

remember to use the chain rule

OpenStudy (ray10):

I'm not sure if I got the correct f'??

OpenStudy (anonymous):

okay so df/dx=1/2 (x^2-y^3-1) ^(1/2- 1) * (2x)

OpenStudy (anonymous):

simplifying yields df/dx= x*(x^2-y^3-1)^(-1/2)

OpenStudy (anonymous):

now you plug in your values

OpenStudy (ray10):

ohh right! Now I get it, let me just go through that again quickly :)

OpenStudy (ray10):

so I plug in my values into the df/dx ?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

so you have 6 (36 -8 -1)^(-1/2)

OpenStudy (anonymous):

whoops that doesn't look right

OpenStudy (ray10):

hmm I seem to get \[\frac{ 2\sqrt{3} }{ 3 }\] what did you get?

OpenStudy (anonymous):

so when you approximate a value for your f(x,y), using partial derivatives should be comparable to your value f(6.1,1.9) see what you get with your original function when x=6.1 and y=1.9

OpenStudy (ray10):

do you have to use \[fxx \times fyy - (fxy)^{2}\] in the steps to working it out? ?

OpenStudy (anonymous):

so this usually works with functions of one variable, i haven't approximated functional evaluations using partial derivatives in a while and was hoping it would be this simple....hmmm maybe you do need another step b/c logically it doesn't make sense

OpenStudy (anonymous):

sorry this is all i could offer :-( i tried

OpenStudy (ray10):

ohh actually I can see a few steps I need to do, I'll try them and get back to you :) thank you for your insight though :)

OpenStudy (anonymous):

You probably just need to find the equation of the tangent plane at some point close to (x,y) = (6.1,1.9) and then use it to approximate f(6.1,1.9)

OpenStudy (ray10):

how do I go about that? :)

OpenStudy (ray10):

Still confused :S